OpenStudy (anonymous):

Identify the conic section that is represented by the equation x2 3y2 + 6x + 2y - 5 = 0

OpenStudy (anonymous):

x^2 + 6x + 9 +3y^2 + 2y = 5 + 9 (x+3)^2 + 3(y^2 + (2/3)y + 1/9) = 14 + 1/3 (x+3)^2 + 3(y + (1/3))^2 = 14 1/3 So it's an ellipse.